OpenStudy (anonymous):

The diameter of ferris wheel is 185 feet, and one complete revolution takes 14 minutes. Find the linear velocity of the person riding on the wheel. Give your answer in miles per hour

OpenStudy (anonymous):

omega = 2 pi /14 min r=d/2 V = r*omega = (2 pi *185 ft/2) / 14 min =pi*185 ft /14 min convert to mph

OpenStudy (anonymous):

thank you! do you know the formula to convert to mph?

OpenStudy (shane_b):

This may help you do the conversion...it's the other way but the steps are basically the same: http://openstudy.com/study#/updates/506241b8e4b0583d5cd30d91

OpenStudy (anonymous):

Im still slightly confused

OpenStudy (anonymous):

im trying to take it from ft/min to mph

OpenStudy (anonymous):

\[\frac{ feet }{ minute}* \frac{ 1 mile }{ 5280 feet } * \frac{ 60 minutes }{1 hour }\]

OpenStudy (anonymous):

ie multiply by 60/5280

OpenStudy (anonymous):

im confused there because i plugged everything in and I got 11100/73920

OpenStudy (anonymous):

I think you forgot the "PI"

OpenStudy (anonymous):

still didnt work but thank you for trying

OpenStudy (anonymous):

how exactly didn't it work?

OpenStudy (anonymous):

i got .22mph and its supposed to be either .53mph, .49mph, .38mph, .65mph, or .47mph

OpenStudy (anonymous):

it's .47

OpenStudy (anonymous):

multiply 11100/73920 by pi

OpenStudy (anonymous):

ooooooh i got it! thank you!